An evacuated observation balloon after filling with helium gas had an increase in weight equal to 2 g. The volume of the balloon at STP must be?

Grade:12th pass

2 Answers

THANGUTURU NIRGUNA
6 years ago
STP value is 22.4 litres.then there is a balloon weight 2 grams.then we have yo multiply the weight *stp value
22.4*2=44.8litres
Yogita Negi
4 years ago
volume=
moles = volume at STP condition/molar volume
i.e. 2/4 = vol./22.4
= 2/4*22.4
= 11.2 L            ...check out this ans.